1. A frustum of a regular hexagonal pyramid has an upper base edge of 5m and a lower base edge of 8.5m. Its lateral area is 160m2. Determine the slant height of the frustum.

a) 2.53 m
b) 3.95 m
c) 4.25 m
d) 15 m

(with solution pls)
1
Expert's answer
2020-12-07T15:59:46-0500

Consider frustrum of a regular hexagonal pyramid

Let "a_1=" the upper base edge, "a_2=" the lower base edge, and "L=" the slant height.

The lateral area "A_L" of the frustum of a regular pyramid is


"A_L=\dfrac{1}{2}n(a_1+a_2)L"

The lateral area "A_L" of the frustum of a regular hexagonal "(n=6)" pyramid is


"A_L=\dfrac{1}{2}(6)(a_1+a_2)L"

"160=\dfrac{1}{2}(6)(5+8.5)L"

"L=\dfrac{160}{40.5}m\approx3.95 m"

b) 3.95 m